# Learning Frequency-Aware Dynamic Transformers for All-In-One Image Restoration

### v1

@@ Abstract
This work aims to tackle the all-in-one image restoration task, which seeks to handle multiple types of degradation with a single model. The primary challenge is to extract degradation representations from the input degraded images and use them to guide the model’s adaptation to specific degradation types. Recognizing that various degradations affect image content differently across frequency bands, we propose a new all-in-one image restoration approach from a frequency perspective, leveraging advanced vision transformers. Our method consists of two main components: a frequency-aware Degradation prior learning transformer (Dformer) and a degradation-adaptive Restoration transformer (Rformer). The Dformer captures the essential characteristics of various degradations by decomposing inputs into different frequency components. By understanding how degradations affect these frequency components, the Dformer learns robust priors that effectively guide the restoration process. The Rformer then employs a degradation-adaptive self-attention module to selectively focus on the most affected frequency components, guided by the learned degradation representations. Extensive experimental results demonstrate that our approach outperforms the existing methods on four representative restoration tasks, including denoising, deraining, dehazing and deblurring. Additionally, our method offers benefits for handling spatially variant degradations and unseen degradation levels.
@@
All-in-one Image Restoration, Frequency-Aware Learning, Vision Transformers.
@@ I Introduction
Image restoration aims to reconstruct high-quality images from degraded ones affected by issues like noise, blur, resolution loss, and various corruptions. Over time, this field has found extensive applications in diverse real-world scenarios, spanning general visual perception, medical imaging, and satellite imaging. Prevailing image restoration efforts center on the meticulous design of task-specific approaches and have demonstrated promising results in tasks such as denoising [1, 2, 3, 4], deraining [5, 6, 7, 8], and deblurring [9, 10, 11, 12]. Despite their success in specific tasks, these approaches often prove inadequate when faced with changes in the degradation task or its severity. This limitation presents significant challenges to their practical use in real-world situations, especially in complex environments. For instance, self-driving cars may encounter consecutive or simultaneous challenges, such as rainy and hazy weather. Consequently, it becomes imperative to develop more generalized approaches capable of recovering images from a variety of unknown degradation types and levels.
Recent studies, e.g., [13, 14], have tried to handle multiple degradations with a multitask learning framework. This involves processing images with different types of degradation by sharing a common backbone and designing task-specific heads. Despite the success of multitask methods in image restoration, those with shared parameters often face the challenge of task interference and still require degradation prior during testing. To avoid these drawbacks, all-in-one image restoration has been studied recently, pioneered by Li et al. [15]. This task aims to address various degradation tasks within a single model. Within the all-in-one framework, the crucial problem to be tackled is how to obtain the degradation representations from the degraded images and how to use the acquired degradation representations in the restoration network. In this work, we propose a new approach to tackle the aforementioned challenges by leveraging advanced vision transformers and recognizing that different degradations impact image content uniquely across frequency bands. Our method comprises two key components: a frequency-aware Degradation Estimation Transformer (Dformer) and a Degradation-Adaptive Restoration Transformer (Rformer).
@@
Rformer functions as a restoration network. The key challenge in designing such a network lies in developing a dynamic module that adapts to various degradation tasks using guidance from degradation representations. Establishing the correlation between the dynamic module and degradation representation is particularly challenging. Li et al. [15] argue that different degradation tasks necessitate different receptive fields within the restoration network. They designed a dynamic module to adjust the receptive field based on the degradation representation. Park et al. [17] introduced an adaptive discriminative filter-based model to explicitly disentangle the restoration network for multiple degradations. Potlapalli et al. [18] proposed a prompt interaction module to enable dynamic interaction between input features and degradation prompts for guided restoration. In contrast, we recognize that different degradation tasks require the restoration model to focus on distinct frequency components of the degraded image. Rformer adapts to these tasks by employing a degradation-adaptive self-attention mechanism, which allows it to adaptively focus on the most affected frequency components, leading to enhanced restoration performance.
To validate the effectiveness of Dformer and Rformer, we conduct extensive experiments. The results demonstrate that our approach surpasses existing methods across four representative restoration tasks: denoising, deraining, dehazing, and deblurring. Furthermore, our method excels in handling spatially variant degradations and previously unseen degradation levels, highlighting its versatility and robustness.
@@ II Related Works
II-A Multiple degradations image restoration
Numerous restoration methods have been developed for specific tasks, utilizing convolutional neural networks [19, 1, 2, 5, 6, 9, 10] or vision transformers [20, 21, 22, 23, 24, 25, 26]. However, these approaches often struggle to generalize beyond particular types and severities of image degradation. To address this limitation, multi-task and all-in-one methods have been proposed, aiming to handle a wider range of degradation types and levels more effectively.
Multi-task methods [13, 14] focus on training a single model to address multiple image restoration tasks simultaneously by incorporating separate modules for each task in parallel at the input and output layers. For example, Chen et al. [13] developed distinct heads and tails for various tasks, with only the backbone being shared among them. Li et al. [14] introduced a task-specific feature extractor to extract common clean features for different adverse weather conditions. However, these methods still rely on specific degradation priors and are unable to handle unknown degradations.
All-in-one methods [15, 18, 27, 28, 29] aim to address a broad spectrum of image restoration tasks with a single, unified model. Unlike multi-task methods, these approaches eliminate the need for prior knowledge of specific degradations or task-specific designs, making them more versatile and efficient in handling various types of image degradation. Wei et al. [30] and Li et al. [15] pioneered this approach by introducing a new method that utilizes contrastive learning to extract degradation representations, thereby guiding the restoration process. Potlapalli et al. [18] proposed a universal and efficient plugin module that employs adjustable prompts to encode degradation-specific information without prior information on the degradations. Park et al. [27] introduced an adaptive discriminant filter-based degradation classifier to explicitly disentangle the network for multiple degradations.
@@ II-A Multiple degradations image restoration
II-B Frequency-aware image restoration
Numerous approaches have emerged to address low-level vision problems, with a focus on frequency analysis. Frequency domain frameworks [31, 32, 33, 34, 35, 36] aim to bridge frequency gaps between sharp and degraded images. For instance, Yang et al. [32] use discrete wavelet transforms to facilitate edge feature extraction. Mao et al. [33] distinguish between blurry and sharp images by processing low- and high-frequency components separately using Fast Fourier Transform. Cui et al. [34] propose a selective frequency module that dynamically separates feature maps into distinct frequency components with learnable filters.
Recent studies [37, 38, 39] have explored biases in frequency domain modules. For instance, the self-attention mechanism in transformers acts as a low-pass filter, while CNN convolutions behave like high-pass filters. This underscores the importance of frequency separation to mitigate model biases by handling different frequencies separately. Our study examines varying frequency objectives across image restoration tasks. Denoising and deraining focus on suppressing high-frequency noise, whereas dehazing and deblurring restore high-frequency details. By addressing inherent frequency biases in transformers’ self-attention modules, we propose a frequency-aware all-in-one image restoration method.
@@ III Method
In this section, we present a new all-in-one image restoration method from a frequency perspective, leveraging advanced vision transformers. Our method comprises two main components: a frequency-aware degradation representation learning transformer (Dformer) and a degradation-adaptive Restoration transformer (Rformer). The Dformer captures the essential characteristics of various degradations by decomposing inputs into different frequency components. By understanding how degradations affect these frequency components, Dformer learns robust priors that effectively guide the restoration process. The Rformer employs a degradation-adaptive self-attention module to adaptively focus on the most affected frequency bands, guided by the acquired degradation representations. This adaptive focus is crucial, as different types of degradations impact image content at various frequency bands.
Formally, given an RGB degraded image $\boldsymbol{I}\in\mathbb{R}^{3\times H\times W}$, its degradation representation $d$ can be obtained using $d=\Phi_{D}(\boldsymbol{I})$ where $\Phi_{D}$ denotes the Dformer. Then the retorted image $\hat{\boldsymbol{I}}$ is obtained with $\hat{\boldsymbol{I}}=\Phi_{R}(\boldsymbol{I},d)$ where $\Phi_{R}$ represents the Rformer. In the following sections, we elaborate on the architectures and optimizations of Dformer $\Phi_{D}$ and Rformer $\Phi_{R}$. The overview of the proposed approach is illustrated in Fig. 1 (a).
@@
Figure 1: Overview of the proposed methods. Dformer learns degradation representation and guides Rformer to achieve all-in-one restoration. Input Frequency Decomposition module utilizes DFT and IDFT processes to decompose the input image into multiple frequency-band images. Input Projection module employs a convolution layer to project the input images into the feature maps. Frequency-Aware Transformer Blocks (FA-TB) is detailed in (b). Output Projection module includes 2D average pooling and two-layer MLP to refine and project degradation representation. Degradation Projection includes a two-layer MLP. The architecture of Rformer follows Uformer, but employs a new degradation-adaptive self-attention mechanism (DA-SA), as detailed in (c), to adaptively handle varying levels and types of image degradation.
III-A Frequency-aware degradation representation learning transformer
We propose Dformer, a frequency-aware transformer specifically designed to learn degradation representations by accounting for the differences in how various degradation types affect image content across frequency bands. Dformer constructs a hierarchical encoder network following the architecture of the Swin Transformer [40], as illustrated in Fig. 1 (a). Dformer incorporates two key designs: 1) Input Frequency Decomposition Module decomposes the input degraded image into distinct frequency bands. 2) Frequency-aware Swin Transformer Block performs self-attention both within and between these frequency bands, effectively learning degradation representations.
@@ III-A Frequency-aware degradation representation learning transformer
This module generates $L$ new images $\{I_{1},I_{2},\ldots,I_{L}\}$, each corresponding to a different frequency band of the degraded image $I$. These images are then passed through a shared $3\times 3$ convolutional layer to extract low-level features. The extracted features are subsequently processed through $K$ shared encoder stages. Each stage consists of $N$ frequency-aware Swin Transformer blocks and a downsampling layer, except for the last stage. After the $K$ encoder stages, we use an output projection, which includes 2D average pooling and two-layer MLP, to generate a degradation representation vector. Next, we will detail the design of our frequency-aware Swin Transformer blocks, which are specifically tailored to capture degradation representations for fully considering every frequency band.
Frequency-aware Transformer block. The widely used Swin Transformer block employs a shifted window-based self-attention mechanism to efficiently capture both local and global contextual information. Unlike the original Swin Transformer block, which processes a single input image $I$, our enhanced block processes $L$ input images, $\{I_{1},I_{2},\ldots,I_{L}\}$, derived from the input frequency decomposition module. To enable the Swin Transformer block to handle multiple frequency-band inputs and fully leverage their contents, we introduce a new frequency-aware transformer block, as illustrated in Fig. 1 (b). This block incorporates new designs in self-attention mechanisms, positional encoding, and masking techniques.
We introduce Intra- and Inter-Band shifted window-based self-attention mechanisms to facilitate adaptive interactions within and between frequency bands. Intra-band self-attention facilitates interactions among distinct pixels within each frequency band, essentially performing self-attention computations independently for each band within the Swin Transformer block. This method ensures complete isolation between different frequency bands, focusing exclusively on intra-band interactions. On the other hand, inter-band self-attention explicitly manages interactions across different frequency bands. Utilizing a window-based strategy, it computes self-attention between pixels from different frequency bands within the same spatial window. This approach allows for a more detailed examination of frequency disparities within localized regions.
@@
III-B Degradation-adaptive restoration transformer
After obtaining the degradation representations, we incorporate them into a restoration transformer (Rformer), as illustrated in Fig. 1 (a). The architecture of Rformer follows Uformer [23], but employs a new degradation-adaptive self-attention mechanism to adaptively focus on the most affected frequency bands, guided by the acquired degradation representations. We illustrate the degradation-adaptive self-attention mechanism in Fig. 1 (c).
Let $\boldsymbol{z}$ be the self-attention map in the transformer block. The frequency band of $\boldsymbol{z}$ can be obtained by Eq. 1. $\operatorname{F-Band}_{k}(\boldsymbol{z})$ denotes the $k$-th frequency band partitioned from the attention maps. After frequency decomposition, the frequency scaling is performed as follows:

### v2

# Learning Dual Transformers for All-In-One Image Restoration from a Frequency Perspective

@@ Abstract
This work aims to tackle the all-in-one image restoration task, which seeks to handle multiple types of degradation with a single model. The primary challenge is to extract degradation representations from the input degraded images and use them to guide the model’s adaptation to specific degradation types. Building on the insight that various degradations affect image content differently across frequency bands, we propose a new dual-transformer approach comprising two components: a frequency-aware Degradation estimation transformer (Dformer) and a degradation-adaptive Restoration transformer (Rformer). The Dformer captures the essential characteristics of various degradations by decomposing the input into different frequency components. By understanding how degradations affect these frequency components, the Dformer learns robust priors that effectively guide the restoration process. The Rformer then employs a degradation-adaptive self-attention module to selectively focus on the most affected frequency components, guided by the learned degradation representations. Extensive experimental results demonstrate that our approach outperforms existing methods in five representative restoration tasks, including denoising, deraining, dehazing, deblurring, and low-light enhancement. Additionally, our method offers benefits for handling, real-world degradations, spatially variant degradations, and unseen degradation levels.
@@
All-in-one Image Restoration, Frequency-Aware Learning, Vision Transformers, Degradation Estimation.
@@ I Introduction
Image restoration aims to reconstruct high-quality images from degraded ones affected by issues like noise, blur, resolution loss, and various corruptions. Over time, this field has found extensive applications in diverse real-world scenarios, spanning general visual perception, medical imaging, and satellite imaging. Prevailing image restoration efforts center on the meticulous design of task-specific approaches and have demonstrated promising results in tasks such as denoising [1, 2, 3, 4], deraining [5, 6, 7, 8], and deblurring [9, 10, 11, 12]. Despite their success in specific tasks, these approaches often prove inadequate when faced with changes in the degradation task or its severity. This limitation presents significant challenges to their practical use in real-world situations, especially in complex environments. For example, self-driving cars may encounter consecutive or simultaneous challenges, such as rainy and hazy weather. Consequently, it becomes imperative to develop more generalized approaches capable of recovering images from a variety of unknown degradation types and levels.
Recent studies, e.g., [13, 14], have tried to handle multiple degradations with a multitask learning framework. This involves processing images with different types of degradation by sharing a common backbone and designing task-specific heads. Despite the success of multitask methods in image restoration, those with shared parameters often face the challenge of task interference and still require degradation prior during testing. To address these drawbacks, all-in-one image restoration has recently received increasing attention [15, 16, 17, 18, 19, 20, 21, 22]. This task aims to address various degradation tasks within a single model. The key challenges in this framework lie in effectively extracting degradation representations from degraded images and utilizing these representations within the restoration network. In this work, we observe through frequency analysis (Section III-A) that different types of degradation affect image content differently across frequency bands. Leveraging this insight, we propose a new dual-transformer approach to tackle these challenges: a Frequency-aware Degradation Estimation Transformer (Dformer) to capture degradation representations and a Degradation-Adaptive Restoration Transformer (Rformer) to utilize these representations for effective image restoration.
@@
Rformer functions as a restoration network. The key challenge in designing such a network lies in developing a dynamic module that adapts to various degradation tasks using guidance from degradation representations. Establishing the correlation between the dynamic module and degradation representation is particularly challenging. Li et al. [15] argue that different degradation tasks necessitate different receptive fields within the restoration network. They designed a dynamic module to adjust the receptive field based on the degradation representation. Park et al. [24] introduced an adaptive discriminative filter-based model to explicitly disentangle the restoration network for multiple degradations. Potlapalli et al. [17] proposed a prompt interaction module to enable dynamic interaction between input features and degradation prompts for guided restoration. In contrast, we recognize that different degradation tasks require the restoration model to focus on distinct frequency components of the degraded image. Rformer adapts to these tasks by employing a frequency-aware self-attention mechanism, which allows it to adaptively focus on the most affected frequency components, leading to enhanced restoration performance.
The contributions of this work are summarized as follows:
  • We propose Dformer, a new degradation estimation transformer that captures the essence of various degradations by decomposing features into frequency components. This frequency-aware design enables Dformer to learn robust priors, effectively guiding the restoration process.
  • We propose Rformer, a restoration transformer that adaptively targets the most affected frequency subbands. By prioritizing frequency bands impacted by different degradations, Rformer ensures precise and effective restoration.
  • We conduct extensive experiments to demonstrate that our approach outperforms the existing methods and analyze its advantages of generalization on five representative restoration tasks, including denoising, deraining, dehazing, deblurring and low-light enhancement.
Next, we discuss the related works in Section II, introduce our approach in detail in Section III, and present our various restoration results in Section IV.
@@ II Related Works
II-A Single degradation image restoration
Numerous methods have been developed to restore clean images from their degraded counterparts. Particularly, convolutional neural networks (CNNs) have demonstrated remarkable efficacy in tasks like image denoising [25, 1, 2, 26], dehazing [27, 28, 29, 30], deraining [31, 32, 33, 34], and deblurring [35, 36, 37, 38]. However, with their potent feature representation capabilities, transformer architectures [39] have emerged as a superior alternative to CNNs in image restoration tasks [40, 41, 42, 43, 44, 45, 46]. For instance, by integrating the hierarchical structure of U-Net [47] with the multi-head self-attention mechanism of Transformers, Swin Transformer [48] effectively balances efficiency and global modeling capabilities through windows-based self-attention. Consequently, Swin Transformer has become a popular choice as the backbone for low-level image restoration models [41, 42, 43, 44]. Liang et al. [41] directly employ Swin Transformer for image restoration tasks, augmenting models with various HQ Image Reconstruction layers to cater to different resolution requirements. Similarly, Wang et al. [43] enhance the decoder layers by replacing the Feed-Forward Network (FFN) with a Locally-enhanced FFN and introducing a learnable modulator before the input of Multi-head Self-attention, allowing for flexible adjustment of feature maps.
However, these methods tailored to specific tasks often struggle to generalize beyond particular types and severities of image degradation. To overcome this challenge, various network design-based approaches have been introduced for general image restoration, demonstrating effectiveness across diverse degradation types [49, 42, 50, 51, 52, 53]. Nevertheless, these approaches still rely on training separate models for different datasets and tasks, adhering to a one-by-one paradigm. In contrast, this paper aims to develop an all-in-one image restoration method capable of handling multiple degradations with a single model.
II-B Multiple degradations image restoration
Multi-task Methods. Recent research [13, 14] has focused on training a single model to address multiple image restoration tasks simultaneously by incorporating separate modules for each task in parallel at the input and output layers. For example, Chen et al. [13] developed distinct heads and tails for various tasks, with only the backbone being shared among them. Li et al. [14] introduced a task-specific feature extractor to extract common clean features for different adverse weather conditions. However, these multi-task methods still rely on prior knowledge of specific degradations and are unable to handle unknown degradations.
All-in-one methods. In contrast to multi-task methods, all-in-one methods [15, 16, 17, 18, 19, 20, 21, 22] aim to tackle a broad spectrum of image restoration tasks using a single, unified model, thus eliminating the need for prior knowledge of specific degradations or task-specific designs. Wei et al. [54] and Li et al. [15] pioneered this approach by introducing a novel method that utilizes contrastive learning to extract degradation representations without prior knowledge of the types and levels of degradations, thereby guiding the restoration process. Potlapalli et al. [17] proposed a universal and efficient plugin module that employs adjustable prompts to encode degradation-specific information without prior information on the degradations. Park et al. [20] introduced an adaptive discriminant filter-based degradation classifier to explicitly disentangle the network for multiple degradations. Building on the success of vision-language models [55], recent studies [56, 57, 58] have integrated language instructions into unified image restoration approaches. However, these methods still require specialized text design for optimal performance.
@@ II-B Multiple degradations image restoration
II-C Frequency-aware image restoration
Numerous approaches have emerged to address various low-level vision problems, with a notable emphasis on frequency analysis. Alongside spatial representation learning, frameworks that operate in the frequency domain [59, 60, 61, 62, 63, 64, 65] primarily aim to bridge the frequency gaps between sharp and degraded image pairs. For instance, Yang et al. [60] utilize discrete wavelet transform and inverse discrete wavelet transform to replace down-sampling and up-sampling, facilitating the extraction of edge features. Mao et al. [61] focus on discerning the differences between blurry and sharp images based on low-frequency and high-frequency components, explicitly segregating the processing of these components using Fast Fourier Transform, and ultimately integrating the residual information from both high and low frequencies. Cui et al. [62] introduce a selective frequency module that dynamically separates feature maps into distinct frequency components using theoretically validated filters, enabling the extraction of high-frequency and low-frequency information in a non-explicit and learnable manner. A common strategy among existing methods involves decomposing features into various frequency components through transformation tools such as wavelet transform, Fourier transform, pooling techniques, and conventional filters, and then treating each component individually.
Moreover, several recent studies [66, 67] have investigated the biases of different modules in the frequency domain. For example, it has been observed that the self-attention mechanism in transformer architectures acts as a low-pass filter, while convolution in CNNs behaves like a high-pass filter. This highlights the importance of frequency separation, as it can help mitigate model biases by handling different frequencies separately. In this study, we delve into the varying frequency objectives across different image restoration tasks. Tasks such as denoising and deraining require the suppression of high-frequency noise, while dehazing and deblurring tasks focus on restoring high-frequency details. By adaptively addressing the inherent frequency biases in self-attention modules of transformer-like models, we propose a new frequency-aware all-in-one image restoration method.
(a) Noisy
(b) Rainy
(c) Hazy
(d) Blurry
Figure 1: Frequency analysis of various degradation types. For each degradation type, the images are the discrete normalized spectrogram (left), the degraded image (right top), and the clean image (right bottom). Noisy and rainy images exhibit a reduced proportion of low-frequency components and an increased proportion of high-frequency components compared to their corresponding clean images. In contrast, hazy and blurry images show the opposite trend.
Figure 2: Visualizing the ratios of low-frequency to high-frequency components between the clean and degraded images using 150 samples. The noisy and rainy samples are positioned in the upper-left region, indicating that the degraded images contain more high-frequency components and fewer low-frequency components than their clean counterparts. In contrast, the hazy and blurred samples are found in the lower-right region, reflecting the opposite trend.
@@ III Method
Figure 3: Overview of the proposed methods. Dformer learns degradation representation and guides Rformer to achieve all-in-one restoration. Input Frequency Decomposition module utilizes DFT and IDFT processes to decompose the input image into multiple frequency-band images. Input Projection module employs a convolution layer to project the input image into the feature maps. Frequency-Aware Transformer Block (FA-TB) is detailed in (b). The image showing a down-arrow within a circle denotes the downsampling layer. Output Projection module includes 2D average pooling and two-layer MLP to refine and project degradation representation. Degradation Projection includes a two-layer MLP. The architecture of Rformer follows Uformer [43], but employs a new degradation-adaptive self-attention mechanism as detailed in (c).
In this work, we propose a new all-in-one image restoration approach from a frequency perspective. We begin by analyzing how various degradation types exhibit distinct characteristics in the frequency domain. Building on these insights, we then introduce the details of our proposed method.
III-A Frequency analysis of various degradation types
In this analysis, we examine four common degradation types that are extensively studied in all-in-one image restoration: noise, rain, haze, and blur. Given the degraded images and their corresponding clean images, We generate spectrograms based on the normalized amplitudes of the residual maps, which are obtained by subtracting the degraded images from the clean images in the frequency domain. To enhance visualization, we divide the spectrograms into 20 equal-width bins, each representing a continuous frequency band. As shown in Fig. 1(a)-(d), we observe that noisy and rainy images exhibit a reduced proportion of low-frequency components and an increased proportion of high-frequency components compared to their corresponding clean images. In contrast, hazy and blurred images show the opposite trend, with a higher concentration of low-frequency components and a lower concentration of high-frequency components.
We further sample 150 image pairs from each of the four degradation types and categorize the normalized amplitudes into two frequency bands: low-frequency and high-frequency, using a method similar to the one described above. For each image pair, we then compute the ratios of low-frequency to high-frequency components between the clean and degraded images. As shown in Fig. 2, we observe that the noisy and rainy image pairs are positioned in the upper-left region, indicating that the degraded images contain more high-frequency components and fewer low-frequency components than their clean counterparts. In contrast, the hazy and blurred image pairs are found in the lower-right region, reflecting the opposite trend.
Building on the analysis above, we conclude that noise and rain primarily degrade images in the high-frequency bands, while haze and blur predominantly affect the low-frequency bands. To address this for all-in-one image restoration, we propose a frequency-aware degradation estimation transformer (Dformer), which captures the key characteristics of various degradations by decomposing inputs into distinct frequency components. By understanding the impact of these degradations on different frequency bands, Dformer learns robust priors that effectively guide the restoration process. We then introduce a degradation-adaptive restoration transformer (Rformer), which incorporates a degradation-adaptive self-attention module. This module enables the model to dynamically focus on the most affected frequency bands, using the degradation representations learned by Dformer. In the following sections, we provide detailed explanations of the architectures and optimizations for Dformer and Rformer. An overview of the proposed approach is shown in Fig. 3(a).
III-B Frequency-aware degradation estimation transformer
Dformer constructs a hierarchical encoder network based on the architecture of the Swin Transformer [48], as illustrated in Fig. 3 (a). Dformer incorporates two key designs: 1) Input Frequency Decomposition Module decomposes the input degraded image into distinct frequency bands. 2) Frequency-aware Swin Transformer Block performs self-attention both within and between these frequency bands, effectively learning degradation representations.
@@ III-B Frequency-aware degradation estimation transformer
Frequency-aware Transformer block. The Swin Transformer block employs a shifted window-based self-attention mechanism to efficiently capture both local and global contextual information. Unlike the original Swin Transformer block, which processes a single input image $\boldsymbol{I}$, our enhanced block processes $L$ input images, ${\boldsymbol{I}_{1},\boldsymbol{I}_{2},\ldots,\boldsymbol{I}_{L}}$, derived from the input frequency decomposition module. To enable the Swin Transformer block to handle multiple frequency-band inputs and fully leverage their contents, we introduce a new frequency-aware transformer block. This block incorporates new designs in self-attention mechanisms, positional encoding, and masking techniques.
We introduce Intra- and Inter-Band shifted window-based self-attention mechanisms to facilitate adaptive interactions within and between frequency bands, as illustrated in Fig. 3 (b). Intra-band self-attention facilitates interactions among distinct pixels within each frequency band, essentially performing self-attention computations independently for each band within the Swin Transformer block. This method ensures complete isolation between different frequency bands, focusing exclusively on intra-band interactions. On the other hand, inter-band self-attention explicitly manages interactions across different frequency bands. Utilizing a window-based strategy, it computes self-attention between pixels from different frequency bands within the same spatial window. This approach allows for a more detailed examination of frequency disparities within localized regions.
@@
Degradation estimation. Given the RGB degraded image $\boldsymbol{I}$, the input frequency decomposition module generates $L$ new images $\{\boldsymbol{I}_{1},\boldsymbol{I}_{2},\ldots,\boldsymbol{I}_{L}\}$, each corresponding to a different frequency band of the degraded image $\boldsymbol{I}$. These images $\{\boldsymbol{I}_{1},\boldsymbol{I}_{2},\ldots,\boldsymbol{I}_{L}\}$ are then passed through a shared $3\times 3$ convolutional layer to extract low-level features. The extracted features are subsequently processed through $K$ shared encoder stages. Each stage consists of $N$ frequency-aware Swin Transformer blocks and a downsampling layer, except for the last stage. The downsampling layer reshapes the flattened features into 2D spatial feature maps, doubles the number of channels, and downsamples the feature maps using a $4\times 4$ convolutional layer with a stride of 2. After the $K$ encoder stages, we use an output projection, which includes 2D average pooling and two-layer MLP, to generate a degradation representation vector.
III-C Degradation-adaptive restoration transformer
After obtaining the degradation representations, we incorporate them into a restoration transformer (Rformer), as illustrated in Fig. 3 (a). Rformer is designed as a general framework adaptable to various transformer-based image restoration architectures, but employs a new degradation-adaptive self-attention mechanism to adaptively focus on the most affected frequency bands, guided by the acquired degradation representations. In this work, we adopt the Uformer [43] as a specific instantiation of Rformer.
Degradation adaptive self-attention. Generally, a self-attention map $\boldsymbol{z}$ in Transformer can be computed as:
$\boldsymbol{z}=softmax(\frac{\boldsymbol{X}\boldsymbol{W}_{Q}(\boldsymbol{X}\boldsymbol{W}_{K})^{T}}{\sqrt{d}})$ (3)
where $\boldsymbol{X}\in\mathbb{R}^{n\times d}$ denotes the feature maps, $n$ and $d$ denote the number of tokens and the feature dimension. $\boldsymbol{W}_{K},\boldsymbol{W}_{Q}\in\mathbb{R}^{d\times d}$ are the weight matrices for key and query, respectively. Formally, the frequency band of the attention maps $\boldsymbol{z}$ can be obtained by Eq. 1. $\operatorname{F-Band}_{k}(\boldsymbol{z})$ denotes the $k$-th frequency band partitioned from the attention maps. After frequency decomposition, the frequency is performed as follows: